A ball is kicked horizontally at 5 m/s from a cliff 50 m high. How far from the base of the cliff will the ball strike the ground?
1
Expert's answer
2020-10-12T07:49:13-0400

"h=gt^2/2\to t=\sqrt{2h/g}=\sqrt{2\cdot 50/9.81}\approx3.2(s)"


"s=vt=5\cdot3.2=16(m)". Answer